# Dissecting inertial clustering and sling dynamics in high-Reynolds number particle-laden turbulence

### v1

@@ I Introduction
(derived from Eq. 7 and similar to Eq. 6) using 100 point-particles with the initial condition $v=\sin(x)$ and $z=\cos(x)$, respectively. The motion of the particles is reduced to one dimension and (in contrast to Eq. 1) a constant, zero-valued flow $u=0$ is prescribed. The results for three different values of $\tau_{p}$ are displayed in Fig. 1, with the top- and bottom-row panels showing the time series of solutions to Eqs. 7 and 8, respectively. We observe that light particles (Fig. 1 and 1) readily evolve in time towards $v=0$ and $z=0$, respectively. The solution remains smooth and well-behaved at long integration times and both signals essentially vanish. In the intermediate case ($\tau_{p}=1$), the velocity gradient develops a sharp peak into the negative direction at $x=\pi$, yet remains bounded (Fig. 1). For $\tau_{p}=3$, Eq. 8 is unstable and $z$ blows up to large negative values (Fig. 1). A fold in $x-v$ space develops, the hallmark of the caustic singularity. This is caused by the de-regularizing effect that inertia exerts on the particle dynamics. Clearly, the relaxation term on the right-hand side of Eq. 7 has a dampening effect on the system: it dictates how fast the solution converges to $v=0$. At large $\tau_{p}$, the relaxation is slow and particles travel far enough to overtake one another in due time; a fold in phase space ensues. Caustic singularities can only occur in Lagrangian flows. The Eulerian equation equivalent to Eq. 7, $\partialderivative{u_{p}}{t}+\partialderivative{(u_{p}^{2}/2)}{x}=-\frac{u_{p}}{\tau_{p}}$, on the other hand, corresponds to the inviscid Burgers equation with a source term on the right-hand side and thus can lead to shocks, if not sufficiently suppressed by the relaxation term (or by viscosity), but the weak solution always remains single-valued, whereas the particles in sling events effectively exhibit multivalued local velocities.
@@ III.1 Instantaneous clustering
Figure 3 shows the divergence of the particle velocity field alongside contours of the local instantaneous $Q$ field at an early instance in time, at which particles have not yet acquired a significant amount of memory, and only very few particles paths have crossed. First, notice how the sharp spectral filter removes the smallest scales of the flow field and thereby smoothens the velocity gradients, as well as massively narrows the range that $Q$ attains. The color map in Fig. 3 is clipped at $Q=\pm 1000$, preventing the contours from being washed out by extreme outliers. Nevertheless, intermittent small-scale regions of intense turbulence are observed, where $Q$, in reality, attains values of up to $10^{5}$. This reflects the nature of turbulence, as the smallest scales exhibit the most extreme vorticity and strain, produced by the non-linear self-amplification of the Navier-Stokes equation. Second, notice the excellent visual agreement between the contours of the divergence patterns (Fig. 3 and 3) and $Q$ (Fig. 3 and 3). ’Particle sinks’ with positive values of $\xi$ emerge in regions where $Q$ is large and positive, i.e., where particles are ejected out of vortices. Particles accumulate, on the other hand, where $\xi$ is negative, leading to an increase of the particle density via Eq. 3. These regions coincide with large negative values of $Q$, i.e., where strain overwhelms vorticity. In other words, the instantaneous flow field explains the spatial distribution of the particles well at this early stage, and Maxey’s centrifuge gives qualitatively accurate predictions of clustering.
Figure 4 visualizes the particles in thin slices of thickness $s=\eta$ after long integration time with the contours of local instantaneous $Q$ in the background. The particles shown have reached steady state, both in terms of the density variance $\langle(n-\langle n\rangle)^{2}\rangle$ and its anti-correlation to $Q$. The sharp spectral filter was applied to the flow in the background of the DNS5 case displayed in Fig. 4 in order to remove turbulent eddies with timescales $\tau<\tau_{p}$ from the flow; in the simulation, however, the unfiltered flow was used.
@@
Clearly, particle clustering is observed in all cases. In the DNS1 case with $St=1$, which is known to exhibit the most intense preferential concentration, clusters are thin, curved structures that are elongated along the local flow topology. Indeed, Eq. 5 seems to predict the spatial distribution of the particles qualitatively well in this case, even after long integration times. As expected [17], these patterns match the typical range of preferential concentration, $\ell\sim 10\eta$, with void sizes spanning into the inertial range of turbulence. A prominent large gap in the particle fields (located in the region $(\tfrac{\pi}{2}\leq x\leq\pi,0\leq y\leq\tfrac{\pi}{2}$)) is seen in all unfiltered cases, coinciding with the largest connected region of increased $Q$ of the coarse field (Fig. 4). Similarly to the DNS1 case, the DNS0.2 particles ($St=0.2$) avoid high-$Q$ and populate low-$Q$ regions, respectively. Particle voids are, however, more exclusive to the most intense vortices. In-between clusters, the particles are distributed more uniformly, and small-scale filamentary patterns are observed. The path-history effect produces fractal clusters with a decrease in the fractal dimension as the Stokes number is increased from $St=0.2$ to $St=1$. In the DNS5 case ($St=5$), the particles are too heavy to resonate with the smallest scales of the flow field, which act to randomize the particle fields and impose their granularity. Once the fine scales are removed, however, qualitative agreement between clusters (voids) and low (high) $Q$ regions is observed, as seen in Fig. 4. Clustering agrees best with the most significant coarse-grained flow contours that match the particle timescale, $\tau_{p}$. Note that the background in Fig. 4 was obtained by filtering the flow field explicitly to match $\tau_{p}$, i.e., so that only turbulent structures of timescale $\tau\leq\tau_{p}$ are removed. This provides evidence for the self-similarity of particle clusters, previously reported by Goto and Vassilicos [30]. The authors’ reasoning is based on the premise that turbulence consists of a hierarchy of self-similar eddies across different length scales, each resonating with particles of a specific $\tau_{p}$.
@@
In the filtered cases (Fig. 4 and 4), instantaneous preferential sampling does not explain clustering well. Both cases exhibit significantly elevated particle number density in low and high $Q$ regions. Moreover, the spatial distributions reveal intricate structures with filamentary patterns. The particle density fields are highly intermittent, with sharp edges separating clusters from voids. The sharp edges are seen because the number of particles remains unchanged between the unfiltered and filtered cases, yet the resolution is higher relative to the smallest scales. Filaments are most conspicuous in the FDNS1 case, where particles fill the domain more uniformly; however, they are also seen in FDNS5, where caustics contribute to the formation of highly irregular patterns. Small-scale filamentary clusters that do not coincide with the local patterns of strain or vorticity have been observed before (e.g. in Fig. 1(b) in Ref. 7), yet the strong misalignment of particle density and $Q$ as well as the large length scale at which filaments are seen are surprising. A possible explanation for the discrepancy between the effectiveness of the centrifuge mechanism to explain clustering in the unfiltered and filtered cases is that the filter removes the regions of the most extreme vorticity and strain, inherent to the dissipative scale of turbulence. We are not convinced by this explanation, however, as, in the smallest scales, the particle fields of the unfiltered cases reveal similar patterns that cannot be attributed to the instantaneous flow field. This is shown in the following section.
@@ III.2 Scale similarity
Scale similarity between the DNS1 and FDNS5 as well as between the DNS0.2 and FDNS1 cases is visually evident, i.e., similar patterns are observed when magnified to the appropriate level. The DNS1 clusters are revealed to have internal structure which where previously (Fig. 4) not visible. These patterns are indicative of multiplicative amplification, which, in simulations of statistical models[7], is observed at length scales even smaller than $\eta$. Numerical studies comparable to the present work usually do not resolve such structures and thus obtain particle fields that can be explained well at sufficiently large scales based on the instantaneous flow fields, not unlike panels 4 or 4 in Fig. 4. Some examples are Fig. 2(a) in Ref. 17, Fig. 2 in Ref. 11, or Fig. 5 in Ref. 31. Based on the visual evidence provided by Fig. 5, however, it must be expected that small-scale filaments are present throughout, yet usually not visible due to the level of magnification or unresolved due to an insufficient number of particles. These filaments are the result of multiplicative amplification, and their arrangement does not align well with the instantaneous flow field.
@@ III.3 Sling events
Patterns of caustics are observed in all cases, as seen in Fig. 6, which shows particles scattering in the computational domain after short integration times. We emphasize that modeling the particle motion using an Eulerian continuum formalism (Eq. 2) is clearly not suitable under these circumstances, as significant numbers of particles undergo caustics and therefore attain multivalued velocities at the same location in space and time (see one-dimensional model presented in Section I). In the filtered cases, caustics are conspicuous thanks to their warped manifold structure at comparatively large length scales, populated by large ensembles of particles. Figure 6 visualizes the peculiar patterns that are associated with sling events at large Stokes number. Heavy particles exhibit high relative velocities and form ’pockets’ [32] (triggering collisions in reality), whereas lighter particles have lower memory capacity and adjust to the flow easily, meaning that lower relative velocities are attained during caustics. This explains the absence of pockets in the FDNS1 case, where caustics appear as lines (sheets in 3D). The pockets in the FDNS5 case are framed by sharp lines of high particle number density, which will be shown to exhibit maximal relative velocities (Fig. 7). A 3D visualization of the pocket character of sling events in the FDNS5 case is provided as supplementary material in Fig. 12.
@@
As described in Section II, Eq. 1 is solved in three dimensions for a large ensemble of particles that lag the turbulent carrier gas flow. We find that caustics form readily, and as early as after $t\approx 2\tau_{\eta}$. To provide an example, we visualize the identification of a sling event in the FDNS5 case in Fig. 7. The large panels depict the particles scattering in a thin slice at three instances in time, colored based on the value of $D$. A slender area of increased relative velocity is easily identified. The sling region, initially a thin line in Fig. 7 (implying a sheet in 3D), attracts particles and grows in width to an elongated cusp. This indicates that, independent of Stokes number, sling events occur on thin sheets, and that pockets (as seen in Fig. 6) develop later in time at large enough $St$. Particles with opposing velocities approach each other from the negative and positive $x$-directions and cluster in the region around $x\approx 0.5$. Here, the particles are focused into a narrow region and the particle number density increases locally by a factor of around $2.5$ (not shown). The transition from a smooth, unambiguous to a multivalued particle velocity field is evident from the insets in Fig. 7, showing the phase space composed of the first velocity and position components of the particles inside the small volume spanning across the caustic and marked by the black rectangle. The signal folds over itself, not unlike the one-dimensional model problem seen in Fig. 1. The insets also reveal that the largest relative velocities exist towards the edge of the caustic manifold, where ’new’, freshly supplied particles are introduced into the sling area. In this region, particle collisions would occur most violently, as particles undergo a sharp increase in relative velocity.
[…]
@@
In Fig. 7, we have illustrated the identification of a sling event by conditioning particles on the nearest-neighbor averaged relative velocity, $D$, and demonstrated how phase-space folds occur in these regions. To build on this, we now depict the flow conditions experienced by particles leading up to the occurrence of a caustic. For the first time to the best of our knowledge, we present 3D renderings of a sling event in Fig. 8. The visualizations are derived from the FDNS1 case, with individual particles represented as points at three consecutive time instances, colored in the locally experienced value of $R$. As described above, the ensemble of sling particles is selected at time $t_{c}=0.6$ using the condition $D(t=t_{c})>D_{C}$. The time instance $t=t_{c}$ is shown in panel 8, clearly revealing that the sling event occurs on a thin, warped sheet in 3D space. The selected ensemble is then tracked backward in time, and the local flow conditions are interpolated to the particle positions to obtain Fig. 8 and 8 at times $t=0.2$ and $t=0.4$, respectively. An additional 3D visualization of the backtracking of sling particle trajectories is provided in the appendix (Fig. 12).
Long before the caustic occurs, the ’epicenter’ of the sling event can be identified in Fig. 8. This area features elevated $R$, but particles have not clustered or collided yet. In the following, the particles inside a dumbbell-like volume approach the epicenter, in which, under the compressive straining of the flow, particles converge on a thin, compact sheet. At time $t=t_{c}-0.2$ (Fig. 8), the ’sling sheet’ commences to take shape, and it is likely that particles in this region already experience elevated relative velocities. Eventually, all particles coalesce into a smooth manifold whose thickness is of the order of the particle diameter (Fig. 8). The fact that we see the sling event taking place on thin sheets is due to the flow topology in the fourth quadrant of the $QR$-plane, where straining acts compressive in one and expanding in the other two directions. Owing to the low effective Stokes number $St_{\mathrm{eff}}=0.2$, however, the particles collide with comparatively low relative velocities, and $D$ decreases rapidly after $t=t_{c}$, as the particles relax to the local flow field and coalesce back into a smooth flow of particles. In the FDNS5 case, on the other hand, particles collide at much higher relative velocities and the sheet structure is not retained later in time (see also Fig. 12). The reason is that the heavier particles cannot adapt to changes of the flow velocity swiftly. This results in higher relative velocities at the time of the sling event and longer times during which they are retained. The crossing of particle paths is responsible for the development of pockets.
@@
While the 3D renderings of Fig. 8 only depicts particles that actually participate in the sling event, Fig. 9 provides a 2D view of particles for which no particular selection was applied. The figure shows particles of the FDNS5 case at time $t=t_{c}+0.2$ (same time instance as in Fig. 7 and 7) as arrows pointing in the direction of the $x,y$-velocities. The contours $Q$ are shown in the background and delineate two vortices, separated by a region of intense strain around $x=0.5$. The flow directions of the particles visualize the sling event that takes place in the $Q<0$ region. Here, the particle velocities are multivalued due to the occurrence of the caustic, as observed in Fig. 7.
@@
Qualitatively, the shapes of the excursions away from $Q=0$ and $R=0$ exhibit only minimal differences between DNS0.2 and DNS1. In fact, both Fig. 10 and 10 enjoy significant visual agreement to Fig. 3 in Ref. 19, which was produced by tracking particles with a Gaussian model at $St=0.3$ and $Ku=22$; as well as to the optimal fluctuations derived in the persistent limit. This indicates that qualitative agreement reaches far into the non-persistent (recall that in the DNS1 case, $St=1$ and $Ku=10.3$) regime and to fully-resolved, high-Reynolds number turbulence. Comparison of Fig. 10 to 10 and Fig. 10 to 10, respectively, verify that the conditions in terms of the fluctuation of $\mathbb{A}$ required to initiate a caustic relax with increasing Stokes number. In other words, heavier particles are driven into sling events more easily, consistent with our observation of the regularizing effect of the $1/\tau_{p}$ term on the right-hand sides of Eqs. 7 and 8. The most probable fluctuations, however, remain similar in shape.
@@
As mentioned earlier, our method cannot predict the exact point in time at which the blow up of $\mathrm{tr}(\mathbb{Z})$ occurs. The vertical lines in Fig. 10 merely depict the point in time at which $D$ is measured compared to the threshold $D_{c}$. Our method’s uncertainty in matching the point in time at which the caustic occurs is reflected by the artifacts of particles undergoing secondary fluctuations visible in Fig. 10. These indicate that the sling event takes place before or after we record the relative velocities. The vast majority of particles, however, are selected at a reasonable time. The variance is lower in the unfiltered cases, as the fluid timescales are much shorter, whereby particles peak in relative velocity during a very brief period of time.
Conspicuously, the sling particles of the DNS5 case (Fig. 10) encounter extreme gradients from early on. Almost all of them are initially located in regions of negative $Q$, with the deviation from $R=0$ lagging slightly but still increasing early in the process. In other words, this leads one to suspect that for a particle with $St=5$, the only prerequisite for participating in a sling event is to be initialized in a region of large negative $Q$ at the start of the simulation and then being ejected into a $R\gg 0$ region shortly thereafter. In truth, however, a history of FVGs that propagates along the Vieillefosse line is required for a caustic to occur, irrespective of the Stokes number, as shown in Fig. 11. At $St=5$, sling events take place on shorter timescales, yet are caused by turbulent flow topologies in a manner very similar to the lower $St$ cases. We do, however, expect that sling events can only occur below a certain Stokes number, as particles become too heavy to respond to the flow at some point and instead move ballistically.

### v3

@@ I Introduction
(derived from Eq. 7 and similar to Eq. 6) using 100 point-particles with the initial condition $v=\sin(x)$ and $z=\cos(x)$, respectively. The motion of the particles is reduced to one dimension and (in contrast to Eq. 1) a constant, zero-valued flow $u=0$ is prescribed. The results for three different values of $\tau_{p}$ are displayed in Fig. 1, with the top- and bottom-row panels showing the time series of solutions to Eqs. 7 and 8, respectively. We observe that light particles (Figs. 1(a,d)) readily evolve in time towards $v=0$ and $z=0$, respectively. The solution remains smooth and well-behaved at long integration times and both signals essentially vanish. In the intermediate case ($\tau_{p}=1$), the velocity gradient develops a sharp peak into the negative direction at $x=\pi$, yet remains bounded (Fig. 1(e). For $\tau_{p}=3$, Eq. 8 is unstable and $z$ blows up to large negative values (Fig. 1(f)). A fold in $x-v$ space develops, the hallmark of the caustic singularity. This is caused by the de-regularizing effect that inertia exerts on the particle dynamics. Clearly, the relaxation term on the right-hand side of Eq. 7 has a dampening effect on the system: it dictates how fast the solution converges to $v=0$. At large $\tau_{p}$, the relaxation is slow and particles travel far enough to overtake one another in due time; a fold in phase space ensues. Caustic singularities can only occur in Lagrangian flows. The Eulerian equation equivalent to Eq. 7, $\partialderivative{u_{p}}{t}+\partialderivative{(u_{p}^{2}/2)}{x}=-\frac{u_{p}}{\tau_{p}}$, on the other hand, corresponds to the inviscid Burgers equation with a source term on the right-hand side and thus can lead to shocks, if not sufficiently suppressed by the relaxation term (or by viscosity), but the weak solution always remains single-valued, whereas the particles in sling events effectively exhibit multivalued local velocities.
@@ III.1 Instantaneous clustering
Figure 3 shows the divergence of the particle velocity field alongside contours of the local instantaneous $Q$ field at an early instance in time, at which particles have not yet acquired a significant amount of memory, and only very few particles paths have crossed. First, notice how the sharp spectral filter removes the smallest scales of the flow field and thereby smoothens the velocity gradients, as well as massively narrows the range that $Q$ attains. The color map in Fig. 3(b) is clipped at $Q=\pm 1000$, preventing the contours from being washed out by extreme outliers. Nevertheless, intermittent small-scale regions of intense turbulence are observed, where $Q$, in reality, attains values of up to $10^{5}$. This reflects the nature of turbulence, as the smallest scales exhibit the most extreme vorticity and strain, produced by the non-linear self-amplification of the Navier-Stokes equation. Second, notice the excellent visual agreement between the contours of the divergence patterns (Figs.3(a,c)) and $Q$ (Figs.3(b,d)). ’Particle sinks’ with positive values of $\xi$ emerge in regions where $Q$ is large and positive, i.e., where particles are ejected out of vortices. Particles accumulate, on the other hand, where $\xi$ is negative, leading to an increase of the particle density via Eq. 3. These regions coincide with large negative values of $Q$, i.e., where strain overwhelms vorticity. In other words, the instantaneous flow field explains the spatial distribution of the particles well at this early stage, and Maxey’s centrifuge gives qualitatively accurate predictions of clustering.
Figure 4 visualizes the particles in thin slices of thickness $s=\eta$ after long integration time with the contours of local instantaneous $Q$ in the background. The particles shown have reached steady state, both in terms of the density variance $\langle(n-\langle n\rangle)^{2}\rangle$ and its anti-correlation to $Q$. The sharp spectral filter was applied to the flow in the background of the DNS5 case displayed in Fig. 4(c) in order to remove turbulent eddies with timescales $\tau<\tau_{p}$ from the flow; in the simulation, however, the unfiltered flow was used.
@@
Clearly, particle clustering is observed in all cases. In the DNS1 case with $St=1$, which is known to exhibit the most intense preferential concentration, clusters are thin, curved structures that are elongated along the local flow topology. Indeed, Eq. 5 seems to predict the spatial distribution of the particles qualitatively well in this case, even after long integration times. As expected [17], these patterns match the typical range of preferential concentration, $\ell\sim 10\eta$, with void sizes spanning into the inertial range of turbulence. A prominent large gap in the particle fields (located in the region $(\tfrac{\pi}{2}\leq x\leq\pi,0\leq y\leq\tfrac{\pi}{2}$)) is seen in all unfiltered cases, coinciding with the largest connected region of increased $Q$ of the coarse field (Fig. 4(c)). Similarly to the DNS1 case, the DNS0.2 particles ($St=0.2$) avoid high-$Q$ and populate low-$Q$ regions, respectively. Particle voids are, however, more exclusive to the most intense vortices. In-between clusters, the particles are distributed more uniformly, and small-scale filamentary patterns are observed. The path-history effect produces fractal clusters with a decrease in the fractal dimension as the Stokes number is increased from $St=0.2$ to $St=1$. In the DNS5 case ($St=5$), the particles are too heavy to resonate with the smallest scales of the flow field, which act to randomize the particle fields and impose their granularity. Once the fine scales are removed, however, qualitative agreement between clusters (voids) and low (high) $Q$ regions is observed, as seen in Fig. 4(c). Clustering agrees best with the most significant coarse-grained flow contours that match the particle timescale, $\tau_{p}$. Note that the background in Fig. 4(c) was obtained by filtering the flow field explicitly to match $\tau_{p}$, i.e., so that only turbulent structures of timescale $\tau\leq\tau_{p}$ are removed. This provides evidence for the self-similarity of particle clusters, previously reported by Goto and Vassilicos [30]. The authors’ reasoning is based on the premise that turbulence consists of a hierarchy of self-similar eddies across different length scales, each resonating with particles of a specific $\tau_{p}$.
@@
In the filtered cases (Figs. 4(d,e)), instantaneous preferential sampling does not explain clustering well. Both cases exhibit significantly elevated particle number density in low and high $Q$ regions. Moreover, the spatial distributions reveal intricate structures with filamentary patterns. The particle density fields are highly intermittent, with sharp edges separating clusters from voids. The sharp edges are seen because the number of particles remains unchanged between the unfiltered and filtered cases, yet the resolution is higher relative to the smallest scales. Filaments are most conspicuous in the FDNS1 case, where particles fill the domain more uniformly; however, they are also seen in FDNS5, where caustics contribute to the formation of highly irregular patterns. Small-scale filamentary clusters that do not coincide with the local patterns of strain or vorticity have been observed before (e.g. in Fig. 1(b) in Ref. 7), yet the strong misalignment of particle density and $Q$ as well as the large length scale at which filaments are seen are surprising. A possible explanation for the discrepancy between the effectiveness of the centrifuge mechanism to explain clustering in the unfiltered and filtered cases is that the filter removes the regions of the most extreme vorticity and strain, inherent to the dissipative scale of turbulence. We are not convinced by this explanation, however, as, in the smallest scales, the particle fields of the unfiltered cases reveal similar patterns that cannot be attributed to the instantaneous flow field. This is shown in the following section.
@@ III.2 Scale similarity
Scale similarity between the DNS1 and FDNS5 as well as between the DNS0.2 and FDNS1 cases is visually evident, i.e., similar patterns are observed when magnified to the appropriate level. The DNS1 clusters are revealed to have internal structure which where previously (Fig. 4(b)) not visible. These patterns are indicative of multiplicative amplification, which, in simulations of statistical models[7], is observed at length scales even smaller than $\eta$. Numerical studies comparable to the present work usually do not resolve such structures and thus obtain particle fields that can be explained well at sufficiently large scales based on the instantaneous flow fields, not unlike panels (a) or (b) in Fig. 4. Some examples are Fig. 2(a) in Ref. 17, Fig. 2 in Ref. 11, or Fig. 5 in Ref. 31. Based on the visual evidence provided by 5(a), however, it must be expected that small-scale filaments are present throughout, yet usually not visible due to the level of magnification or unresolved due to an insufficient number of particles. These filaments are the result of multiplicative amplification, and their arrangement does not align well with the instantaneous flow field.
@@ III.3 Sling events
Patterns of caustics are observed in all cases, as seen in Fig. 6, which shows particles scattering in the computational domain after short integration times. We emphasize that modeling the particle motion using an Eulerian continuum formalism (Eq. 2) is clearly not suitable under these circumstances, as significant numbers of particles undergo caustics and therefore attain multivalued velocities at the same location in space and time (see one-dimensional model presented in Section I). In the filtered cases, caustics are conspicuous thanks to their warped manifold structure at comparatively large length scales, populated by large ensembles of particles. Figure 6(d) visualizes the peculiar patterns that are associated with sling events at large Stokes number. Heavy particles exhibit high relative velocities and form ’pockets’ [32] (triggering collisions in reality), whereas lighter particles have lower memory capacity and adjust to the flow easily, meaning that lower relative velocities are attained during caustics. This explains the absence of pockets in the FDNS1 case, where caustics appear as lines (sheets in 3D). The pockets in the FDNS5 case are framed by sharp lines of high particle number density, which will be shown to exhibit maximal relative velocities (Fig. 7). A 3D visualization of the pocket character of sling events in the FDNS5 case is provided as supplementary material in Fig. 12(b).
@@
As described in Section II, Eq. 1 is solved in three dimensions for a large ensemble of particles that lag the turbulent carrier gas flow. We find that caustics form readily, and as early as after $t\approx 2\tau_{\eta}$. To provide an example, we visualize the identification of a sling event in the FDNS5 case in Fig. 7. The large panels depict the particles scattering in a thin slice at three instances in time, colored based on the value of $D$. A slender area of increased relative velocity is easily identified. The sling region, initially a thin line in Fig. 7 (a) (implying a sheet in 3D), attracts particles and grows in width to an elongated cusp. This indicates that, independent of Stokes number, sling events occur on thin sheets, and that pockets (as seen in Fig. 6(d)) develop later in time at large enough $St$. Particles with opposing velocities approach each other from the negative and positive $x$-directions and cluster in the region around $x\approx 0.5$. Here, the particles are focused into a narrow region and the particle number density increases locally by a factor of around $2.5$ (not shown). The transition from a smooth, unambiguous to a multivalued particle velocity field is evident from the insets in Fig. 7, showing the phase space composed of the first velocity and position components of the particles inside the small volume spanning across the caustic and marked by the black rectangle. The signal folds over itself, not unlike the one-dimensional model problem seen in Fig. 1. The insets also reveal that the largest relative velocities exist towards the edge of the caustic manifold, where ’new’, freshly supplied particles are introduced into the sling area. In this region, particle collisions would occur most violently, as particles undergo a sharp increase in relative velocity.
@@
In Fig. 7, we have illustrated the identification of a sling event by conditioning particles on the nearest-neighbor averaged relative velocity, $D$, and demonstrated how phase-space folds occur in these regions. To build on this, we now depict the flow conditions experienced by particles leading up to the occurrence of a caustic. For the first time to the best of our knowledge, we present 3D renderings of a sling event in Fig. 8. The visualizations are derived from the FDNS1 case, with individual particles represented as points at three consecutive time instances, colored in the locally experienced value of $R$. As described above, the ensemble of sling particles is selected at time $t_{c}=0.6$ using the condition $D(t=t_{c})>D_{C}$. The time instance $t=t_{c}$ is shown in panel (c), clearly revealing that the sling event occurs on a thin, warped sheet in 3D space. The selected ensemble is then tracked backward in time, and the local flow conditions are interpolated to the particle positions to obtain Figs. 8(a,b) at times $t=0.2$ and $t=0.4$, respectively. An additional 3D visualization of the backtracking of sling particle trajectories is provided in the appendix (Fig. 12).
Long before the caustic occurs, the ’epicenter’ of the sling event can be identified in Fig. 8(a). This area features elevated $R$, but particles have not clustered or collided yet. In the following, the particles inside a dumbbell-like volume approach the epicenter, in which, under the compressive straining of the flow, particles converge on a thin, compact sheet. At time $t=t_{c}-0.2$ (Fig. 8(b)), the ’sling sheet’ commences to take shape, and it is likely that particles in this region already experience elevated relative velocities. Eventually, all particles coalesce into a smooth manifold whose thickness is of the order of the particle diameter (Fig. 8(c)). The fact that we see the sling event taking place on thin sheets is due to the flow topology in the fourth quadrant of the $QR$-plane, where straining acts compressive in one and expanding in the other two directions. Owing to the low effective Stokes number $St_{\mathrm{eff}}=0.2$, however, the particles collide with comparatively low relative velocities, and $D$ decreases rapidly after $t=t_{c}$, as the particles relax to the local flow field and coalesce back into a smooth flow of particles. In the FDNS5 case, on the other hand, particles collide at much higher relative velocities and the sheet structure is not retained later in time (see also Fig. 12(b)). The reason is that the heavier particles cannot adapt to changes of the flow velocity swiftly. This results in higher relative velocities at the time of the sling event and longer times during which they are retained. The crossing of particle paths is responsible for the development of pockets.
@@
While the 3D renderings of Fig. 8 only depicts particles that actually participate in the sling event, Fig. 9 provides a 2D view of particles for which no particular selection was applied. The figure shows particles of the FDNS5 case at time $t=t_{c}+0.2$ (same time instance as in Figs. 7(c,d)) as arrows pointing in the direction of the $x,y$-velocities. The contours $Q$ are shown in the background and delineate two vortices, separated by a region of intense strain around $x=0.5$. The flow directions of the particles visualize the sling event that takes place in the $Q<0$ region. Here, the particle velocities are multivalued due to the occurrence of the caustic, as observed in Fig. 7(d).
@@
Qualitatively, the shapes of the excursions away from $Q=0$ and $R=0$ exhibit only minimal differences between DNS0.2 and DNS1. In fact, both Figs. 10(a,b) enjoy significant visual agreement to Fig. 3 in Ref. 19, which was produced by tracking particles with a Gaussian model at $St=0.3$ and $Ku=22$; as well as to the optimal fluctuations derived in the persistent limit. This indicates that qualitative agreement reaches far into the non-persistent (recall that in the DNS1 case, $St=1$ and $Ku=10.3$) regime and to fully-resolved, high-Reynolds number turbulence. Comparison of Figs. 10(a) to (b) and Figs. 10(d) to (e), respectively, verify that the conditions in terms of the fluctuation of $\mathbb{A}$ required to initiate a caustic relax with increasing Stokes number. In other words, heavier particles are driven into sling events more easily, consistent with our observation of the regularizing effect of the $1/\tau_{p}$ term on the right-hand sides of Eqs. 7 and 8. The most probable fluctuations, however, remain similar in shape.
@@
As mentioned earlier, our method cannot predict the exact point in time at which the blow up of $\mathrm{tr}(\mathbb{Z})$ occurs. The vertical lines in Fig. 10 merely depict the point in time at which $D$ is measured compared to the threshold $D_{c}$. Our method’s uncertainty in matching the point in time at which the caustic occurs is reflected by the artifacts of particles undergoing secondary fluctuations visible in Fig. 10(d). These indicate that the sling event takes place before or after we record the relative velocities. The vast majority of particles, however, are selected at a reasonable time. The variance is lower in the unfiltered cases, as the fluid timescales are much shorter, whereby particles peak in relative velocity during a very brief period of time.
Conspicuously, the sling particles of the DNS5 case (Fig. 10(c)) encounter extreme gradients from early on. Almost all of them are initially located in regions of negative $Q$, with the deviation from $R=0$ lagging slightly but still increasing early in the process. In other words, this leads one to suspect that for a particle with $St=5$, the only prerequisite for participating in a sling event is to be initialized in a region of large negative $Q$ at the start of the simulation and then being ejected into a $R\gg 0$ region shortly thereafter. In truth, however, a history of FVGs that propagates along the Vieillefosse line is required for a caustic to occur, irrespective of the Stokes number, as shown in Fig. 11. At $St=5$, sling events take place on shorter timescales, yet are caused by turbulent flow topologies in a manner very similar to the lower $St$ cases. We do, however, expect that sling events can only occur below a certain Stokes number, as particles become too heavy to respond to the flow at some point and instead move ballistically.